OpenStudy (anonymous):

Find the arc length L of a circle with a radius of 6 feet and an arc measure of 120°. Give the answer in terms of π.

OpenStudy (anonymous):

@mathmale , @wolf1728

OpenStudy (anonymous):

here arc measure of 120 = 120π /180 = 2π/3 radian hence length of the arc is given by 6* 2π/3= 4π

OpenStudy (anonymous):

So it's 4π?

OpenStudy (anonymous):

This is one of my three hardest final questions. I've been working on it for literally 8 hours.

OpenStudy (anonymous):

length of arc = radius * (arc measure of the angle)

OpenStudy (anonymous):

6*120=720

OpenStudy (anonymous):

angle shoul be in radians

OpenStudy (anonymous):

use 180° = π radians (by d way the questio is asking to leave the answer in terms of π )

OpenStudy (anonymous):

The circumference is proportional to the arc lenth.

OpenStudy (anonymous):

\[ \frac{x}{2\pi r } =\frac{120^\circ}{360^{\circ}} \]

OpenStudy (anonymous):

Do I cross multiply?

OpenStudy (anonymous):

Nah, just multiply by \(2\pi r\) and simplify.

OpenStudy (anonymous):

Multiply what by 2πr? (it's been a long day sorry)

OpenStudy (anonymous):

Both sides.

OpenStudy (anonymous):

We are solving for \(x\), which is the arc length.

OpenStudy (anonymous):

I'm a bit lost.

OpenStudy (anonymous):

So 2πr*x/2πr?

OpenStudy (anonymous):

Wait, have you not learned any algebra?

OpenStudy (anonymous):

No I have. I've just got a case of the stupid right now.

OpenStudy (anonymous):

What you get is: \[ x = \frac{120}{360}\cdot 2\pi r \]

OpenStudy (anonymous):

.33+2π=2.09

OpenStudy (anonymous):

* times

OpenStudy (anonymous):

in this case, \(r\) is 4

OpenStudy (anonymous):

How'd you get 4?

OpenStudy (anonymous):

I mean 6

OpenStudy (anonymous):

Lol, how'd you get 6?

OpenStudy (anonymous):

circle with a radius of 6 feet

OpenStudy (anonymous):

Oh wait yes.

OpenStudy (anonymous):

So my answer is 12.5?

OpenStudy (anonymous):

Exact answer is \(4\pi\), but it evaluates to about 12.57

OpenStudy (anonymous):

Okay awesome. Thank you so much!